OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft he can have 1000 earthworms. How mant worms can he have if his square compost has a side length that is 8 times longer? B) 64,000 C) 256,000 D) 32,000

OpenStudy (anonymous):

set up the ratio and solve for x. 1000 worms /16 square feet is what you know. the new box is 32*32 (4*8 becuase the new box has 8 times the lenght of the sides of the orinigial box) \[\frac{ 1000 }{ 16 }=\frac{ x }{ 1024}\] 1024 being 32*32. Solve for x = 1000*1024/16

OpenStudy (anonymous):

thx so much!

OpenStudy (anonymous):

64,000